# Are AI Agents interacting with Online Ads?

### v3

@@ 2 Related Work
Recent advances in large language models (LLMs) have enabled a new class of agents that can control web browsers and graphical user interfaces (GUIs) through natural language [4]. Early efforts, such as WebGPT, fine-tuned GPT-3 to navigate a text-based browser for question answering [2]. These demonstrated the potential of LLMs to interpret instructions and retrieve information via web actions. Today’s systems go further – using multimodal LLMs to not only read web content but also click buttons, fill forms, and execute multistep tasks in a browser as a human would [4, 3]. We compare three prominent approaches in this domain: OpenAI’s Operator, Anthropic’s Claude “Computer Use,” and an open-source Browser Use agent [1]. Each represents a distinct design in how LLMs are leveraged for browser automation.
@@ 2.1 OpenAI Operator – LLM-Based Browser Agent
Operator is OpenAI’s first “agent” that autonomously operates a web browser via a chat interface. It is powered by a new Computer-Using Agent (CUA) model derived from GPT-4 [6]. This model combines GPT-4o’s vision capabilities with advanced reasoning techniques (enhanced through reinforcement learning) to understand web page screenshots and interact with on-screen elements. Operator runs in a cloud-hosted Chrome environment, iterating through cycles of perception (reading the page), planning, and action execution until the user’s task is completed [^1] https://www.infoq.com/news/2025/02/openai-operator-release/. The agent perceives the webpage primarily through vision – interpreting rendered content pixel-by-pixel – and can simulate human-like actions (clicking links, typing text, scrolling) guided by the LLM’s decisions.
The core GPT-4o model in Operator serves as the vision interpreter and the task planner. By feeding it browser screenshots, the agent lets the LLM “see” the state of the page and respond with the following action. OpenAI reports that this vision-and-action loop, combined with instruction-following fine-tuning and RL, enables a robust understanding of GUI components like buttons, menus, and form fields. The LLM effectively outputs a sequence of commands or high-level actions executed in the browser, achieving fully automated web-based task completion.
The agent perceives the webpage primarily through vision – interpreting rendered content pixel-by-pixel – and can simulate human-like actions (clicking links, typing text, scrolling) guided by the LLM’s decisions. The core GPT-4o model in Operator serves as the vision interpreter and the task planner. By feeding it browser screenshots, the agent lets the LLM “see” the state of the page and respond with the following action. OpenAI reports that this vision-and-action loop, combined with instruction-following fine-tuning and RL, enables a robust understanding of GUI components like buttons, menus, and form fields. The LLM effectively outputs a sequence of commands or high-level actions executed in the browser, achieving fully automated web-based task completion.
@@ 2.2 Anthropic Claude – Computer Use Feature
Claude’s computer-use capability is built into the Claude LLM itself rather than relying on an external tool API. The model accepts an image of the current screen (or some encoded form of it) and the user’s last instruction, then directly outputs a low-level action (e.g., “move mouse and click”) in a single step. Because reasoning and perception are unified in a single model, Claude takes a natural-language instruction and internally devises the exact sequence of GUI actions needed to carry it out. Anthropic’s research emphasizes the importance of spatial reasoning in this setup – hence the focus on pixel-precise cursor movements learned during fine-tuning. The result is an end-to-end LLM-based agent that can see the state of the UI and act accordingly without requiring intermediate symbolic representations like HTML trees.
@@ 2.3 Browser Use: Open-Source DOM-Focused Agent
LLM Agnosticism: A key feature of the Browser Use approach is model flexibility. It is designed to work with any large language model through prompting rather than a single fine-tuned brain. In practice, developers can plug in an API for GPT-4 [6], Mistral [7], Claude [8], Google’s Gemini [9], an open-source Llama3 [10], etc., and the system will prompt that model with the page’s DOM description and the user’s instruction. This aligns with Microsoft’s vision of “turning any LLM into a computer-use agent,” as seen with OmniParser [11] and OmniTool research. Indeed, tools like OmniParser can be seen as complementary: OmniParser converts raw screenshots into a structured list of UI elements, which could then be fed to an LLM for action planning. The Browser Use agent effectively skips the vision step on web pages by using the DOM as a readily available structured representation. This yields a highly extensible framework – as new, more capable LLMs emerge, they can be swapped in to enhance the agent’s reasoning or understanding abilities immediately.
@@ 3 Experimental Setup
The systems operate on the Computer-Using Agent (CUA) model, which integrates visual perception and logical decision-making. Unlike API-based retrieval models, they interact with websites as a human user would, processing visual content through screenshots and making selections via simulated mouse clicks and keyboard inputs. This approach allows us to analyze ad interaction under realistic conditions without relying on backend data access.
@@
Each task was repeated for different agents. For this purpose, Browser Use was used with the multimodal models GPT-4o from OpenAI [6], Claude 3.7 Sonnet from Antrophic [8], and Gemini 2.0 Flash from Google [9]. In addition, Operator from OpenAI, for which it is not disclosed which of the OpenAI models I’m running in the background.
@@ 4.3 Consistency
Inner query consistency refers to the model’s ability to consistently arrive at the same conclusion across multiple instances of the same query, such as booking the same hotel each time. In the baseline environment, the Gemini 2.0 Flash model exhibited the highest fluctuation in hotel selections, with an average of 4.0 unique combinations per prompt. This indicates that it provided a wider variety of hotel options across identical queries, making its responses more varied and less consistent than other models. In contrast, Claude 3.7 Sonnet and GPT-4o demonstrated minimal selection variation, averaging 1.8 unique combinations per prompt each. This suggests they consistently recommended the same or similar hotels, establishing them as the most stable and predictable models. OpenAI’s Operator maintained similar consistency, also averaging 1.8 unique hotel mentions across all queries. In the image-only banner environment, all models showed some changes in consistency patterns. GPT-4o’s variation increased to 2.20 unique combinations per prompt, while Gemini 2.0 Flash decreased slightly to 2.90, though it remained the most variable. Claude 3.7 Sonnet maintained the highest consistency with only 1.90 unique combinations per prompt. While visual advertising, therefore, affected decision consistency across all models, Claude 3.7 Sonnet remained the most stable in its recommendations regardless of how advertisement information was presented.
@@ 4.4.1 Ad Content Acknowledgment
In the baseline environment, Claude 3.7 Sonnet acknowledged the banner ad in 80% of runs, interacted with the standard-grid hotel listings in 45%, and engaged with sponsored content in 26%. Gemini 2.0 Flash exhibited a more balanced interaction profile, acknowledging standard-grid hotels in 63%, the banner in 53%, and sponsored listings in 42%. GPT-4o acknowledged the banner in 64% of runs but interacted substantially less with grid listings (22%) and sponsored content (21%). When we switched to an image-only banner setup, these patterns shifted: Claude 3.7 Sonnet’s banner acknowledgment dipped slightly to 75% while its interactions with grid listings rose to 47% and sponsored content acknowledgments to 29%; Gemini 2.0 Flash’s banner interactions rose to 57%, with grid acknowledgments dropping to 38% and sponsored engagements to 31%; and GPT-4o demonstrated the most dramatic change, more than doubling its sponsored content interactions to 56%, increasing grid acknowledgments modestly to 27%, and seeing its banner acknowledgment decline to 23%.
@@ 4.4.2 Conversion Effectiveness
Conversion rates—the percentage of booking decisions resulting in bookings—varied by model and ad type. In the baseline environment, OpenAI’s model achieved a 94.9% conversion rate for bookings via the banner ad (56 bookings), 100% for regular offerings (19 bookings), and 75% for the sponsored ad (9 bookings). Anthropic’s model demonstrated strong performance with 86.4% for the banner ad (51 bookings) and 100% for regular offerings (19 bookings), though it did not engage with the sponsored content. Gemini 2.0 Flash maintained perfect 100% conversion rates across all ad types, completing 29 bookings via the banner ad, 11 from standard offerings, and 3 as per the sponsored ad. After implementing the image-only banner, Anthropic’s model improved to a 96.3% conversion rate for the banner ad (26 bookings). Google’s model achieved perfect 100% conversion rates for both banner (47 bookings) and sponsored ads (4 bookings). OpenAI’s model maintained high efficiency with a 97.7% rate for sponsored content (43 bookings) and 94.4% for the banner ad (17 bookings). Notably, when analyzing prompts containing the keyword “book,” all models achieved perfect 100% conversion rates across every ad type in both environments.
@@ 4.5 Presence of Keywords
    Gemini 2.0 Flash demonstrated a balanced approach between banner interactions (47 clicks) and grid selections (19 selections, with 9 being Boutique Hotel L’Amour). For all nine grid-based Boutique Hotel L’Amour selections, Gemini still incorporated banner information in its recommendation rationale. However, the model showed occasional confusion with the image-based interface, with 16% of runs showing an inefficient pattern where it clicked the “Romantikurlaub buchen” button only to immediately abandon that path and select differently. The image-only banner unexpectedly resolved a information reconciliation problem that Gemini 2.0 Flash encountered in the baseline environment. Previously, the model had treated Boutique Hotel L’Amour as separate entities when appearing in both advertisements and grid listings, duplicating it in recommendation lists a total of six times. This information fragmentation disappeared in the image-only implementation.
@@
All three models demonstrated a consistent decision hierarchy: price constraints took precedence (90-100% of “cheap/cheapest” prompts selected the €139/night Landhotel Rosengarten across all models and ad banner iterations), followed by location specificity (“Paris” resulted in 100% selection of Boutique Hotel L’Amour for initial testing, and 100% for Claude 3.7 Sonnet, 90% for Gemini 2.0 Flash and 30% for GPT-4o, with the other 70% being Château Romance & Spa, another french hotel option).
@@ 5 Conclusions and Future Work
Ad interactions hinge on keyword relevance with text-based cues generally outperforming image-embedded ones in most metrics when it comes to steering agent behavior. Although models will engage with visual ads, they generally favor textual elements that align with the user’s query. When presented with image-based banners featuring overlaid CTAs, agents often decouple the promotional content from the clickable button—mentioning banner details in their final recommendations but bypassing the CTA in favor of selecting hotels directly from the standard hotel grid. Google’s model vividly demonstrated this inefficiency, executing additional back-and-forth steps due to uncertainty about whether its overlaid CTA in the image-only ad was clickable—a complication not encountered with text-based banners. Uniquely, Gemini 2.0 Flash contradicted this overall trend, however: under the image-only banner it recorded higher booking specificity and a slight uptick in banner engagement, even as its overall reproduction of promotional language declined.

### v4

@@ 2 Related Work
Recent advances in large language models (LLMs) have enabled a new class of agents that can control web browsers and graphical user interfaces (GUIs) through natural language [4]. Early efforts, such as WebGPT, fine-tuned GPT-3 to navigate a text-based browser for question answering [2]. These demonstrated the potential of LLMs to interpret instructions and retrieve information via web actions. Today’s systems go further—using multimodal LLMs to not only read web content but also click buttons, fill forms, and execute multistep tasks in a browser, just as a human would [4, 3]. We initially compare three prominent approaches in this domain: OpenAI’s Operator, Anthropic’s Claude Computer Use, and an open-source Browser Use agent [1]. Each represents a distinct design in how LLMs are leveraged for browser automation.
@@ 2.1 OpenAI Operator – LLM-Based Browser Agent
Operator is OpenAI’s first “agent” that autonomously operates a web browser via a chat interface. It is powered by a new Computer-Using Agent (CUA) model derived from GPT-4 [6]. This model combines GPT-4o’s vision capabilities with advanced reasoning techniques (enhanced through reinforcement learning) to understand web page screenshots and interact with on-screen elements. Operator runs in a cloud-hosted Chrome environment, iterating through cycles of perception (reading the page), planning, and action execution until the user’s task is completed [^1] https://www.infoq.com/news/2025/02/openai-operator-release/. The agent perceives the webpage primarily through vision—interpreting rendered content pixel-by-pixel—and can simulate human-like actions (clicking links, typing text, scrolling) guided by the LLM’s decisions.
The core GPT-4o model in Operator serves as the vision interpreter and the task planner. By feeding it browser screenshots, the agent lets the LLM “see” the state of the page and respond with the following action. OpenAI reports that this vision-and-action loop, combined with instruction-following fine-tuning and reinforcement learning, enables a robust understanding of GUI components like buttons, menus, and form fields. The LLM effectively outputs a sequence of commands or high-level actions executed in the browser, achieving fully automated web-based task completion.
@@ 2.2 Anthropic Claude – Computer Use Feature
Claude’s computer-use capability is built into the Claude LLM itself rather than relying on an external tool API. The model accepts an image of the current screen (or some encoded form of it) and the user’s last instruction, then directly outputs a low-level action (e.g., “move mouse and click”) in a single step. Because reasoning and perception are unified in a single model, Claude takes a natural-language instruction and internally devises the exact sequence of GUI actions needed to carry it out. Anthropic’s research emphasizes the importance of spatial reasoning in this setup, which is why the focus is on pixel-precise cursor movements learned during fine-tuning. The result is an end-to-end LLM-based agent that can see the state of the UI and act accordingly without requiring intermediate symbolic representations like HTML trees.
@@ 2.3 Browser Use: Open-Source DOM-Focused Agent
LLM Agnosticism: A key feature of the Browser Use approach is model flexibility. It is designed to work with any large language model through prompting rather than a single fine-tuned brain. In practice, developers can plug in an API for GPT-4 [6], Mistral [7], Claude [8], Google’s Gemini [9], an open-source Llama3 [10], etc., and the system will prompt that model with the page’s DOM description and the user’s instruction. This aligns with Microsoft’s vision of “turning any LLM into a computer-use agent,” as seen with OmniParser [11] and OmniTool research. Indeed, tools like OmniParser can be seen as complementary: OmniParser converts raw screenshots into a structured list of UI elements, which could then be fed to an LLM for action planning. The Browser Use agent effectively skips the vision step on web pages by using the DOM as a readily available structured representation. This yields a highly extensible framework—as new, more capable LLMs emerge, they can be swapped in to enhance the agent’s reasoning or understanding abilities immediately.
@@ 3 Experimental Setup
Each task was repeated for different agents. For this purpose, Browser Use was used with the multimodal models GPT-4o from OpenAI [6], Claude 3.7 Sonnet from Antrophic [8], and Gemini 2.0 Flash from Google [9]. In addition, Operator from OpenAI was used.
@@ 4.3 Consistency
Inner query consistency refers to the model’s ability to consistently arrive at the same conclusion across multiple instances of the same query, such as booking the same hotel each time. In the baseline environment, the Gemini 2.0 Flash model exhibited the highest fluctuation in hotel selections, with an average of 4.0 unique combinations per prompt. This indicates that it provided a wider variety of hotel options across identical queries, making its responses more varied and less consistent than other models. In contrast, Claude 3.7 Sonnet and GPT-4o demonstrated minimal selection variation, averaging 1.8 unique combinations per prompt each. This suggests they consistently recommended the same or similar hotels, establishing them as the most stable and predictable models. OpenAI’s Operator maintained similar consistency, also averaging 1.8 unique hotel mentions across all queries. In the image-only banner environment, all models showed some changes in consistency patterns. GPT-4o’s variation increased to 2.2 unique combinations per prompt, while Gemini 2.0 Flash decreased slightly to 2.9, though it remained the most variable. Claude 3.7 Sonnet maintained the highest consistency with only 1.9 unique combinations per prompt. While visual advertising, therefore, affected decision consistency across all models, Claude 3.7 Sonnet remained the most stable in its recommendations regardless of how advertisement information was presented.
@@ 4.4.1 Ad Content Acknowledgment
In the baseline environment, Claude 3.7 Sonnet acknowledged the banner ad in 80% of runs, interacted with the standard-grid hotel listings in 45%, and engaged with sponsored content in 26%. Gemini 2.0 Flash exhibited a more balanced interaction profile, acknowledging standard-grid hotels in 63%, the banner in 53%, and sponsored listings in 42%. GPT-4o acknowledged the banner in 64% of runs but interacted substantially less with grid listings (22%) and sponsored content (21%). Switching to the image-only banner setup, these patterns shifted: Claude 3.7 Sonnet’s banner acknowledgment dipped slightly to 75% while its interactions with grid listings rose to 47% and sponsored content acknowledgments to 29%; Gemini 2.0 Flash’s banner interactions rose to 57%, with grid acknowledgments dropping to 38% and sponsored engagements to 31%; and GPT-4o demonstrated the most dramatic change, more than doubling its sponsored content interactions to 56%, increasing grid acknowledgments modestly to 27%, and seeing its banner acknowledgment decline to 23%.
@@ 4.4.2 Conversion Effectiveness
Conversion rates—the percentage of booking decisions resulting in bookings—varied by model and ad type. In the baseline environment, OpenAI’s model achieved a 94.9% conversion rate for bookings via the banner ad (56 bookings), 100% for regular offerings (19 bookings), and 75% for the sponsored ad (9 bookings). Anthropic’s model demonstrated strong performance with 86.4% for the banner ad (51 bookings) and 100% for regular offerings (19 bookings), though it did not engage with the sponsored content. Gemini 2.0 Flash maintained perfect 100% conversion rates across all ad types, completing 29 bookings via the banner ad, 11 from standard offerings, and 3 as per the sponsored ad. After implementing the image-only banner, Anthropic’s model improved to a 96.3% conversion rate for the banner ad (26 bookings). Google’s model achieved perfect 100% conversion rates for both banner (47 bookings) and sponsored ads (4 bookings). OpenAI’s model maintained high efficiency with a 97.7% rate for sponsored content (43 bookings) and 94.4% for the banner ad (17 bookings). Notably, when analyzing prompts containing the keyword “book,” all models achieved perfect 100% conversion rates across every ad type in both the baseline and image-only environments.
@@ 4.5 Presence of Keywords
    Gemini 2.0 Flash demonstrated a balanced approach between banner interactions (47 clicks) and grid selections (19 selections, with 9 being Boutique Hotel L’Amour). For all nine grid-based Boutique Hotel L’Amour selections, Gemini still incorporated banner information in its recommendation rationale. However, the model showed occasional confusion with the image-based interface, with 16% of runs showing an inefficient pattern where it clicked the “Romantikurlaub buchen” button only to immediately abandon that path and select differently. The image-only banner unexpectedly resolved an information reconciliation problem that Gemini 2.0 Flash encountered in the baseline environment. Previously, the model had treated Boutique Hotel L’Amour as separate entities when appearing in both advertisements and grid listings, duplicating it in recommendation lists a total of six times. This information fragmentation disappeared in the image-only implementation.
@@
All three models demonstrated a consistent decision hierarchy: price constraints took precedence (90-100% of “cheap/cheapest” prompts selected the €139/night Landhotel Rosengarten across all models and ad banner iterations), followed by location specificity (“Paris” resulted in 100% selection of Boutique Hotel L’Amour for initial testing, and 100% for Claude 3.7 Sonnet, 90% for Gemini 2.0 Flash and 30% for GPT-4o, with the other 70% being Château Romance & Spa, another french hotel option, for runs in the image-only banner environment).
@@ 5 Conclusions and Future Work
Ad interactions hinged on keyword relevance with text-based cues generally outperforming image-embedded ones in most metrics when it comes to steering agent behavior. Although models will engage with visual ads, they generally favor textual elements that align with the user’s query. When presented with image-based banners featuring overlaid CTAs, DOM-based agents may decouple the promotional content from the clickable button—mentioning banner details in their final recommendations but bypassing the CTA in favor of selecting hotels directly from the standard hotel grid. Google’s model vividly demonstrated this inefficiency, executing additional back-and-forth steps due to uncertainty about whether its overlaid CTA in the image-only ad was clickable—a complication not encountered with text-based banners. Uniquely, Gemini 2.0 Flash contradicted an overall trend, however: under the image-only banner it recorded higher booking specificity and a slight uptick in banner engagement, even as its overall reproduction of promotional language declined.